# Multigranular Evaluation for Brain Visual Decoding

### v1

@@ Abstract
Existing evaluation protocols for brain visual decoding predominantly rely on coarse metrics that obscure inter-model differences, lack neuroscientific foundation, and fail to capture fine-grained visual distinctions. To address these limitations, we introduce BASIC, a unified, multigranular evaluation framework that jointly quantifies structural fidelity, inferential alignment, and contextual coherence between decoded and ground truth images. For the structural level, we introduce a hierarchical suite of segmentation-based metrics, including foreground, semantic, instance, and component masks, anchored in granularity-aware correspondence across mask structures. For the semantic level, we extract structured scene representations encompassing objects, attributes, and relationships using multimodal large language models, enabling detailed, scalable, and context-rich comparisons with ground-truth stimuli. We benchmark a diverse set of visual decoding methods across multiple stimulus-neuroimaging datasets within this unified evaluation framework. Together, these criteria provide a more discriminative, interpretable, and comprehensive foundation for measuring brain visual decoding methods.
1 Introduction
Recent advances in brain visual decoding takagi2023high ; ozcelik2023brain ; scotti2023reconstructing ; xia2024umbrae have achieved remarkable success in reconstructing visual stimuli from the neural activations. However, the evaluation protocols commonly used in this field remain limited in several critical aspects. First, current metrics often saturate across state-of-the-art models, limiting their discriminative capacity and obscuring substantive differences in decoded results. Second, these metrics often lack a neuroscientific foundation, failing to capture the perceptual validity of decoded outputs and their alignment with human-like perception. Third, prevailing evaluation strategies typically fail to reflect the multilevel and structured nature of visual perception, neglecting key components such as object semantics, scene understanding, and contextual reasoning. These limitations hinder rigorous benchmarking of brain decoding models and obscure the specific dimensions along which reconstructions succeed or fall short.
@@ 1 Introduction
Brain visual decoding reconstructs visual experiences from neural activations – aiming to recover not only the appearance of stimuli but also their structure, semantics, and perceptual salience. A well-decoded reconstruction should reflect what the subject consciously perceived, preserving salient objects, their attributes, spatial configuration, and overall scene coherence. In line with human visual perception, which is shaped by attention, context, and prior knowledge, successful brain visual decoding should align with the hierarchical structure of vision, ranging from low-level pixel appearance to high-level scene semantics. This necessitates that brain visual decoding preserves both perceptual accuracy and semantic integrity, capturing salient objects, their attributes, and spatial context in alignment with the subject’s attention, while maintaining consistent relationships, interactions, and overall scene coherence.
@@
Current evaluation protocols, such as pixel-wise correlation or feature-based similarity, often fall short in capturing the full complexity of brain visual decoding. Low-level metrics tend to overlook scene semantics and perceptual plausibility, while high-level black-box measures conflate multiple alignment aspects into a single score, offering limited diagnostic insight. They struggle to determine whether the details in the “decoded” output truly originates from brain signals, or if they are instead hallucinated constructs based on prototypical co-occurrences from pretrained generative models conditioned on scene or object labels. Moreover, existing metrics are often saturated, assigning uniformly high scores across diverse methods and thus failing to capture fine-grained distinctions in reconstruction quality. We argue that an effective brain visual decoding metric should be neuroscientifically interpretable, grounded in principles of human visual perception, and meet three key desiderata. First, it should be multigranular, capturing perceptual alignment across multiple abstraction levels – from basic object identification and segmentation to semantic and spatial reasoning about attributes and interactions. Second, it should be semantically aligned with human perception, reflecting the way humans interpret scenes, objects, and their relationships in a coherent and meaningful manner. Third, it should be diagnostically informative, offering interpretable feedback on what is correct, what is missing, and where the reconstruction fails. Specifically, it should localize and characterize semantic and structural errors in the decoded outputs, such as object misidentification, incorrect attributes, or implausible interactions.
@@
BASIC offers a comprehensive view of decoding performance across modalities (image, video, 3D) and neuroimaging types (fMRI, EEG). Our framework facilitates both quantitative comparisons and qualitative diagnostics, allowing for fine-grained benchmarking of brain decoding models across datasets. BASIC aims to (1) offer a more detailed and interpretable evaluation of brain decoding results, (2) quantify semantic plausibility in terms more closely aligned with human cognitive processes, and (3) facilitate comparison of decoding methods within a unified evaluation framework applicable across diverse stimulus-neuroimaging datasets. We hope our method contributes to establishing a more systematic foundation for evaluating advancements in visual brain decoding.
2 Related Work
@@
The task of brain visual decoding aims to reconstruct perceived visual stimuli, such as images, videos, or 3D shapes, from recorded neural activations. Recent progress in this domain has been closely tied to advancements in computational modeling frameworks. Early methods primarily involved training neural networks from scratch to learn mappings between brain activity and visual features. However, these models often suffered from limited fidelity and exhibited artifacts and poor visual quality in the reconstructed outputs. Recent developments have led to significant improvements due to the emergence of multimodal generative models radford2021learning ; rombach2022high ; xu2022versatile and large-scale brain-stimulus datasets wen2018neural ; allen2022massive ; gao2024mind3d ; guo2025neuro3d . These resources have facilitated a new generation of decoding paradigms that leverage intermediate representations from pretrained generative models and align them with neural responses through various strategies, including linear regression ozcelik2023brain ; takagi2023high , diffusion priors scotti2023reconstructing ; scotti2024mindeye2 , or feature-wise reconstruction xia2024umbrae ; xia2025mevox . Recent efforts have also focused on eliminating dependence on subject-specific encoders scotti2024mindeye2 ; wang2024mindbridge ; xia2025vindex , demonstrating promising progress toward subject-independent brain decoding. Besides the commonly used fMRI-image Natural Scenes Dataset (NSD) allen2022massive , a growing body of research has explored alternative combinations of neuroimaging modalities and stimuli, such as fMRI-video wen2018neural , fMRI-3D gao2024mind3d , EEG-image grootswagers2022human , EEG-video liu2024eeg2video , and EEG-3D guo2025neuro3d decoding. These explorations broaden the scope of brain decoding and offer new insights into the neural representation of dynamic and immersive visual experiences.
@@
The evaluation metrics for visual brain decoding lack a universally accepted standard. Different stimulus-neuroimaging combinations employ varying evaluation protocols. For instance, the following eight metrics are commonly used on the NSD: PixCorr, SSIM wang2004image , AlexNet krizhevsky2017imagenet -2/5, Inception szegedy2016rethinking , CLIP radford2021learning , EffNet tan2019efficientnet , and SwAV caron2020unsupervised . The first four metrics are considered low-level, focusing on perceptual similarity and pixel-wise or structural correspondence. In contrast, the latter four are used to evaluate semantic decoding or high-level representations, emphasizing the overall theme or content of the reconstructed images. For other stimulus-neuroimaging setups, evaluation protocols remain inconsistent. Metrics such as 2-way or 50-way classification accuracy guo2025neuro3d , CLIP-based Pearson correlation gong2024neuroclips , mask-matching ratios li2025decofuse , and modality-specific evaluation metrics xia2024dream ; gao2024mind3d++ have all been reported in the recent literature. However, prior metrics overlook the hierarchical nature of perception and struggle to distinguish models with subtle differences. In contrast, our BASIC framework offers a unified, multigranular evaluation across structural, inferential, and contextual dimensions, enabling more nuanced and interpretable comparisons.
Figure 1: BASIC evaluates decoded reconstructions along two axes: high-level semantic (BASIC-H, Sec. 3.2) and low-level structural (BASIC-L, Sec. 3.3) similarities. For the semantic axis (inferential and contextual), we extract and compare structured representations from reconstructed and ground-truth images. For the structural axis, we compute mask-based matching across fine-grained segmentation types of identified scenes and objects: salient, semantic, instance, and parts.
3 BASIC: Evaluating Brain Visual Decoding
BASIC provides a versatile evaluation framework for brain visual decoding across diverse combinations of visual stimuli and neuroimaging modalities. This section details the evaluation dimensions (Sec. 3.1) and the two complementary modules of our BASIC metric: BASIC-H (Sec. 3.2) integrates the inferential and contextual dimensions of BASIC into a unified measure of high-level semantic correspondence, while BASIC-L (Sec. 3.3) quantifies low-level structural alignment.
3.1 Evaluation Dimension
To better understand how brain decoding models capture visual semantics across abstraction levels, we design a structured evaluation protocol from conceptual dimensions of human visual perception: scene, object, attributes, and relations. Each is further divided into sub-categories as shown in Tab. 1.
@@ 3.1 Evaluation Dimension
Camera captures the photographer’s perspective and natural conditions during the shot, including the viewpoint (e.g., frontal, top-down), motion (e.g., zoom, pan), and lighting.
These dimensions are motivated by principles in visual neuroscience and cognitive psychology desimone1984stimulus ; puce1996differential ; kanwisher1997fusiform , and aligned with the structure of scene understanding in large multimodal models dong2024benchmarking ; lu2024benchmarking ; liu2023visual . These dimensions are then organically integrated into structural, inferential, and contextual similarity components within our framework, converted into two sub-indicators, as outlined below.
Table 1: Evaluation dimension. Each result is examined from five dimensions, each decomposed into sub-components. This enables a multigranular evaluation consistent with visual perception principles.
Dimension Details
@@
relation spatial relation, part-whole relation, physical interaction, kinematics (relative motion, speed)
@@
3.2 BASIC-H
Detailed description generation.
To support multigranular evaluation, we first prompt MLLMs to produce semantically rich descriptions from images. This description prompt extracts detailed information across key semantic components corresponding to the semantic dimensions outlined in Tab. 1, including scene context, objects, object attributes, inter-object relations, and camera viewpoint (when applicable). Recent studies indicate that state-of-the-art MLLMs can generate captions on par with human experts, albeit with occasional hallucinations lu2024benchmarking ; dong2024benchmarking .
Semantic graph construction.
Given detailed captions, we first segment them into individual sentences using NLTK loper2002nltk . We then extract key visual elements – objects, attributes, and relations – using a T5-based factual parser li2023factual . To handle objects mentioned across multiple sentences in the caption, we consolidate references to the same object into subcaptions and bind attributes to the correct instances, avoiding indiscriminate merging. For directional relations (e.g., subject-object pairs), we concatenate relevant sentences to maintain relational coherence.
Structured semantic matching.
The matching process aligns the extracted objects, attributes, and relations in three steps: (a) exact matching, which identifies direct lexical matches between visual elements from the reconstruction and the reference stimuli; (b) synonym matching, which aligns semantically equivalent terms, such as “building” and “edifice” for objects; “bright blue” and “azure” for attributes; “next to” and “beside” for relations; and (c) semantic matching, which computes cosine similarity for any remaining unmatched elements based on contextual meaning. This multi-step process ensures that all visual elements, whether directly mentioned or linguistically varied, are appropriately matched across the reconstructed and reference images.
Finally, we compute semantic alignment scores based on the matching results. To mitigate the effects of omissions and hallucinations in MLLM outputs, we evaluate precision, recall, and F1 scores for each type of visual element (object, attribute, relation), based on the matching outcomes. The overall BASIC-H score is computed as a weighted sum of the three F1 scores, with weights of $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ assigned to objects, attributes, and relations, respectively. Precision penalizes hallucinations by measuring the proportion of correctly matched elements among all generated ones, whereas recall captures omissions by evaluating how completely the reference content is recovered. Their combination in F1 score enables a balanced assessment of both over- and under-generation issues common in MLLM outputs. Further details and examples of BASIC-H are provided in Sec. C.1.
3.3 BASIC-L
Semantic instance categorization.
We prompt MLLMs to generate structured semantic annotations from images. These annotations include object-centric identifications that specify spatial roles (e.g., foreground or background), semantic categories (e.g., “dog”, “tree”), and part-level component decompositions (e.g., “head”, “leg”). This process incorporates a multi-level decomposition of visual elements, capturing a hierarchical view of the scene components and facilitating further grounded segmentation and multigranular structural alignment. This categorization prompt for layered instance annotations and corresponding examples are provided in Appendix Tab. S7 and Tab. S8, respectively.
Progressive granular segmentation.
The categories from these structured annotations are then used as prompts for referring expression comprehension methods ren2024grounded ; kirillov2023segment , which produces segmentation masks for each identified object and component. The segmentation results are influenced by both text and box threshold settings, where only the highest-similarity boxes exceeding the box threshold and words with similarity scores above the text threshold are considered as predicted labels liu2024grounding .
Hierarchical score calculation.
We compare these predicted masks against those from the reference image across four segmentation granularities: salient (foreground categories, F), binary (foreground and background categories, B), semantic (all distinct categories, S), instance (individual instances of all identified object categories, I), and part (sub-object components, mostly salient foreground objects, P), computing intersection-over-union (IoU) and average precision (AP) for each.
We aggregate saliency, semantic, instance, and part-level segmentation scores into the overall BASIC-L metric via a weighted sum of respective IoUs using weights of $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $\beta_{4}$ for each granularity. This scheme prioritizes global layout and object-level coherence while accounting for fine-grained details—reflecting the limited granularity of information captured during the brain data acquisition experiments. Further implementation details and examples of BASIC-L are provided in Sec. C.2.
@@ 3.3 BASIC-L
NSD allen2022massive
SDRecon takagi2023high 55.59 53.05 53.79 10.12 38.73 14.96 40.15 38.71 39.06 35.31
BrainDiffuser ozcelik2023brain 57.87 59.11 58.09 13.35 45.82 19.43 43.20 44.42 43.50 39.71
MindEye scotti2023reconstructing 62.94 60.64 61.26 18.14 51.17 25.06 49.98 48.42 48.84 44.30
DREAM xia2024dream 65.63 63.06 63.56 18.97 50.68 25.92 53.45 53.21 52.91 46.37
MindEye2 scotti2024mindeye2 62.57 62.12 61.72 17.86 50.16 24.71 49.72 49.17 49.07 44.39
MindBridge wang2024mindbridge 59.00 58.35 58.19 13.70 45.69 19.49 45.75 45.78 45.43 40.16
UMBRAE xia2024umbrae 62.00 61.86 61.44 17.51 51.32 24.49 48.91 48.71 48.45 44.06
NeuroPictor huo2024neuropictor 63.00 61.05 61.38 17.92 50.13 24.66 49.62 49.08 48.98 44.21
NeuroVLA shen2024neuro 65.36 65.03 64.57 21.27 53.73 28.65 53.80 52.86 52.95 47.88
SepBrain wang2024unibrain 62.10 60.19 60.57 16.62 48.71 23.31 48.03 47.55 47.44 43.04
UniBrain wang2024unibrain 59.03 58.21 58.07 13.27 43.89 19.02 45.55 45.58 45.25 39.89
STTM liu2025see 64.50 62.50 62.88 19.53 51.70 26.64 51.17 50.28 50.36 45.88
MindTuner gong2025mindtuner 62.55 62.64 61.95 18.06 49.18 24.73 50.22 50.10 49.80 44.63
BrainGuard tian2025brainguard 63.63 62.36 62.43 18.66 52.40 25.84 51.25 50.72 50.60 45.43
EEG-Things  grootswagers2022human
ATM li2024visual 44.54 44.51 43.77 9.65 38.48 14.46 36.74 36.63 36.28 30.55
CognitionCapturer zhang2025cognitioncapturer 42.20 43.99 42.59 7.56 34.56 11.67 34.31 35.74 34.70 28.64
CC2017 wen2018neural
MinD-Video chen2023cinematic 47.56 45.05 45.34 6.95 29.27 10.57 32.42 31.10 31.32 28.63
NeuroClips gong2024neuroclips 62.82 61.68 61.28 17.12 50.82 24.30 53.91 56.33 54.42 45.12
DecoFuse li2025decofuse 47.35 47.12 46.74 6.26 28.12 9.77 32.75 32.18 32.11 29.03
SEED-DV liu2024eeg2video
EEG2Video liu2024eeg2video 66.05 65.75 64.36 25.43 51.85 32.20 54.63 56.02 54.59 49.54
fMRI-Shape gao2024mind3d
MinD-3D gao2024mind3d 41.52 36.42 37.59 13.52 44.00 19.15 42.49 41.12 41.26 30.95
MinD-3D++ gao2024mind3d++ 41.35 37.92 38.12 20.03 61.89 27.80 44.82 43.39 43.57 35.08
EEG-3D guo2025neuro3d
Neuro-3D guo2025neuro3d 35.79 32.86 34.26 6.32 27.92 10.32 23.41 29.87 26.24 23.08
Table 3: BASIC-L scores across datasets. This metric evaluates the low-level structural correspondence between reconstructed and reference images at four granularities: foreground saliency, semantic consistency, instance separation, and part-level delineation. Higher is better.

### v2

@@ Abstract
Existing evaluation protocols for brain visual decoding predominantly rely on coarse metrics that obscure inter-model differences, lack neuroscientific foundation, and fail to capture fine-grained visual distinctions. To address these limitations, we introduce BASIC, a unified, multigranular evaluation framework that jointly quantifies structural fidelity, inferential alignment, and contextual coherence between decoded and ground-truth images. For the structural level, we introduce a hierarchical suite of segmentation-based metrics, including foreground, semantic, instance, and component masks, anchored in granularity-aware correspondence across mask structures. For the semantic level, we extract structured scene representations encompassing objects, attributes, and relationships using multimodal large language models, enabling detailed, scalable, and context-rich comparisons with ground-truth stimuli. We benchmark a diverse set of visual decoding methods across multiple stimulus-neuroimaging datasets within this unified evaluation framework. Together, these criteria provide a more discriminative, interpretable, and comprehensive foundation for evaluating brain visual decoding methods.
Codehttps://github.com/weihaox/BASIC
Introduction
Recent advances in brain visual decoding (Takagi and Nishimoto 2023a; Ozcelik and VanRullen 2023; Scotti et al. 2023; Xia et al. 2024a) have achieved remarkable success in reconstructing visual stimuli from the neural activations. However, the evaluation protocols commonly used in this field remain limited in several critical aspects. First, current metrics often saturate across state-of-the-art models, limiting their discriminative capacity and obscuring substantive differences in decoded results. Second, these metrics often lack a neuroscientific foundation, failing to capture the perceptual validity of decoded outputs and their alignment with human-like perception. Third, prevailing evaluation strategies typically fail to reflect the multilevel and structured nature of visual perception, neglecting key components such as object semantics, scene understanding, and contextual reasoning. These limitations hinder rigorous benchmarking of brain decoding models and obscure the specific dimensions along which reconstructions succeed or fall short.
@@ Introduction
Brain visual decoding aims to reconstruct visual experiences from neural activations, recovering not only the appearance of stimuli but also their structure, semantics, and perceptual salience. A well-decoded reconstruction should reflect what the subject consciously perceived, preserving salient objects, their attributes, spatial configuration, and overall scene coherence. Since human visual perception is shaped by attention, context, and prior knowledge, brain decoding must align with the hierarchical nature of vision, spanning from low-level pixel patterns to high-level semantic understanding. Therefore, effective brain visual decoding requires both perceptual accuracy and semantic integrity. It should faithfully capture salient elements and spatial context in line with the subject’s attention, while maintaining consistent inter-object relationships and scene-level coherence.
@@
Current evaluation protocols, such as pixel-wise correlation or feature-based similarity, often fall short in capturing the full complexity of brain visual decoding. Low-level metrics tend to overlook scene semantics and perceptual plausibility, while high-level black-box measures conflate multiple alignment aspects into a single score, offering limited diagnostic insight. They struggle to determine whether the “decoded” details truly originates from brain signals, or if they are instead hallucinated constructs based on prototypical co-occurrences from pretrained generative models conditioned on scene or object labels. Moreover, existing metrics are often saturated, assigning uniformly high scores across diverse methods and thus failing to capture fine-grained distinctions in reconstruction quality. We argue that an effective brain visual decoding metric should be neuroscientifically interpretable, grounded in principles of human visual perception, and meet three key desiderata. First, it should be multigranular, capturing perceptual alignment across multiple abstraction levels – from basic object identification and segmentation to semantic and spatial reasoning about attributes and interactions. Second, it should be semantically aligned with human perception, reflecting the way humans interpret scenes, objects, and their relationships in a coherent and meaningful manner. Third, it should be diagnostically informative, offering interpretable feedback on what is correct, what is missing, and where the reconstruction fails. Specifically, it should localize and characterize semantic and structural errors in decoded outputs, such as object misidentification, incorrect attributes, or implausible interactions.
@@
BASIC offers a comprehensive view of decoding performance across modalities (image, video, 3D) and neuroimaging types (fMRI, EEG). Our framework facilitates both quantitative comparisons and qualitative diagnostics, allowing for fine-grained benchmarking of brain decoding models across datasets. BASIC aims to (1) offer a more detailed and interpretable evaluation of brain decoding results, (2) quantify semantic plausibility in terms more closely aligned with human cognitive processes, and (3) facilitate comparison of decoding methods within a unified evaluation framework applicable across diverse stimulus-neuroimaging datasets. We hope our method contributes to establishing a more systematic foundation for brain visual decoding evaluation.
Figure 1: BASIC evaluates decoded reconstructions along two axes: high-level semantic (BASIC-H) and low-level structural (BASIC-L) similarities. For the semantic axis (inferential and contextual), we extract and compare structured representations from reconstructed and ground-truth images. For the structural axis, we compute mask-based matching across fine-grained segmentation types of identified scenes and objects: salient, semantic, instance, and parts.
Related Work
@@
The task of brain visual decoding aims to reconstruct perceived visual stimuli, such as images, videos, or 3D shapes, from recorded neural activations. Recent progress in this domain has been closely tied to advancements in computational modeling frameworks. Early methods primarily involved training neural networks from scratch to learn mappings between brain activity and visual features. However, these models often suffered from limited fidelity and exhibited artifacts and poor visual quality in the reconstructed outputs. Recent developments have led to significant improvements due to the emergence of multimodal generative models (Radford et al. 2021; Rombach et al. 2022; Xu et al. 2023) and large-scale brain-stimulus datasets (Wen et al. 2018; Allen et al. 2022; Gao et al. 2024; Guo et al. 2025). These resources have facilitated a new generation of decoding paradigms that leverage intermediate representations from pretrained generative models and align them with neural responses through various strategies, including linear regression (Ozcelik and VanRullen 2023; Takagi and Nishimoto 2023a), diffusion priors (Scotti et al. 2023, 2024), or feature-wise reconstruction (Xia et al. 2024a; Xia and Öztireli 2025b, a). Recent efforts have also focused on eliminating dependence on subject-specific encoders (Scotti et al. 2024; Wang et al. 2024a; Xia et al. 2024a; Tian et al. 2025; Gong et al. 2025), demonstrating promising progress toward subject-independent brain decoding. Besides the commonly used fMRI-image Natural Scenes Dataset (NSD) (Allen et al. 2022), a growing body of research has explored alternative combinations of neuroimaging modalities and stimuli, such as fMRI-video (Wen et al. 2018), fMRI-3D (Gao et al. 2024), EEG-image (Grootswagers et al. 2022), EEG-video (Liu et al. 2024b), and EEG-3D (Guo et al. 2025) decoding. These explorations broaden the scope of brain decoding and offer new insights into the neural representation of dynamic and immersive visual experiences.
@@
The evaluation metrics for visual brain decoding lack a universally accepted standard. Different stimulus-neuroimaging combinations employ varying evaluation protocols. For instance, the following eight metrics are commonly used on NSD: PixCorr, SSIM (Wang et al. 2004), AlexNet (Krizhevsky, Sutskever, and Hinton 2017)-2/5, Inception (Szegedy et al. 2016), CLIP (Radford et al. 2021), EffNet (Tan and Le 2019), and SwAV (Caron et al. 2020). The first four metrics are considered low-level, focusing on perceptual similarity and pixel-wise or structural correspondence. In contrast, the latter four are used to evaluate semantic decoding or high-level representations, emphasizing the overall theme or content of the reconstructed images. For other stimulus-neuroimaging setups, evaluation protocols remain inconsistent. Metrics such as $n$-way classification accuracy (Guo et al. 2025), CLIP-based Pearson correlation (Gong et al. 2024), mask-matching ratios (Li et al. 2025), and modality-specific evaluation metrics (Gao et al. 2025) have all been reported in the recent literature. However, prior metrics overlook the hierarchical nature of perception and struggle to distinguish models with subtle differences. In contrast, our BASIC framework offers a unified, multigranular evaluation across structural, inferential, and contextual dimensions, enabling more nuanced and interpretable comparisons.
BASIC: Evaluating Brain Visual Decoding
BASIC provides a versatile evaluation framework for brain visual decoding across diverse combinations of visual stimuli and neuroimaging modalities. This section details the evaluation dimensions and the two complementary modules of our BASIC metric: BASIC-H integrates the inferential and contextual dimensions of BASIC into a unified measure of high-level semantic correspondence, while BASIC-L quantifies low-level structural alignment.
Evaluation Dimension
To gain deeper insight into how brain decoding models represent visual semantics across different levels of abstraction, we develop a structured evaluation protocol grounded in key conceptual dimensions of human visual perception: scenes, objects, attributes, and relations. Each dimension is further divided into subcategories as outlined in Table 1.
@@
Camera captures the photographer’s perspective and natural conditions during the shot, including viewpoint (e.g., frontal, top-down), motion (e.g., zoom, pan), and lighting.
These dimensions are motivated by principles in visual neuroscience and cognitive psychology (Desimone et al. 1984; Puce et al. 1996), and aligned with the structure of scene understanding in large multimodal models (Dong et al. 2024; Lu et al. 2025; Liu et al. 2023). These dimensions are then organically integrated into structural, inferential, and contextual similarity components within our framework, converted into two sub-indicators, as outlined below.
Table 1: Evaluation dimension and details.
Dims. Details
@@
relation spatial or part-whole relation, interaction, kinematics (motion, speed)
@@
NSD (Allen et al. 2022)
SDRecon (Takagi and Nishimoto 2023a) 55.59 53.05 53.79 10.12 38.73 14.96 40.15 38.71 39.06 35.31
BrainDiffuser (Ozcelik and VanRullen 2023) 57.87 59.11 58.09 13.35 45.82 19.43 43.20 44.42 43.50 39.71
MindEye (Scotti et al. 2023) 62.94 60.64 61.26 18.14 51.17 25.06 49.98 48.42 48.84 44.30
DREAM (Xia et al. 2024b) 65.63 63.06 63.56 18.97 50.68 25.92 53.45 53.21 52.91 46.37
MindEye2 (Scotti et al. 2024) 62.57 62.12 61.72 17.86 50.16 24.71 49.72 49.17 49.07 44.39
MindBridge (Wang et al. 2024a) 59.00 58.35 58.19 13.70 45.69 19.49 45.75 45.78 45.43 40.16
UMBRAE (Xia et al. 2024a) 62.00 61.86 61.44 17.51 51.32 24.49 48.91 48.71 48.45 44.06
NeuroPictor (Huo et al. 2024) 63.00 61.05 61.38 17.92 50.13 24.66 49.62 49.08 48.98 44.21
NeuroVLA (Shen et al. 2024) 65.36 65.03 64.57 21.27 53.73 28.65 53.80 52.86 52.95 47.88
SepBrain (Wang et al. 2024b) 62.10 60.19 60.57 16.62 48.71 23.31 48.03 47.55 47.44 43.04
UniBrain (Wang et al. 2024b) 59.03 58.21 58.07 13.27 43.89 19.02 45.55 45.58 45.25 39.89
STTM (Liu et al. 2025) 64.50 62.50 62.88 19.53 51.70 26.64 51.17 50.28 50.36 45.88
MindTuner (Gong et al. 2025) 62.55 62.64 61.95 18.06 49.18 24.73 50.22 50.10 49.80 44.63
BrainGuard (Tian et al. 2025) 63.63 62.36 62.43 18.66 52.40 25.84 51.25 50.72 50.60 45.43
EEG-Things  (Grootswagers et al. 2022)
ATM (Li et al. 2024) 44.54 44.51 43.77 9.65 38.48 14.46 36.74 36.63 36.28 30.55
CognitionCapturer (Zhang et al. 2025) 42.20 43.99 42.59 7.56 34.56 11.67 34.31 35.74 34.70 28.64
CC2017 (Wen et al. 2018)
MinD-Video (Chen, Qing, and Zhou 2023) 47.56 45.05 45.34 6.95 29.27 10.57 32.42 31.10 31.32 28.63
NeuroClips (Gong et al. 2024) 62.82 61.68 61.28 17.12 50.82 24.30 53.91 56.33 54.42 45.12
DecoFuse (Li et al. 2025) 47.35 47.12 46.74 6.26 28.12 9.77 32.75 32.18 32.11 29.03
SEED-DV (Liu et al. 2024b)
EEG2Video (Liu et al. 2024b) 66.05 65.75 64.36 25.43 51.85 32.20 54.63 56.02 54.59 49.54
fMRI-Shape (Gao et al. 2024)
MinD-3D (Gao et al. 2024) 41.52 36.42 37.59 13.52 44.00 19.15 42.49 41.12 41.26 30.95
MinD-3D++ (Gao et al. 2025) 41.35 37.92 38.12 20.03 61.89 27.80 44.82 43.39 43.57 35.08
EEG-3D (Guo et al. 2025)
Neuro-3D (Guo et al. 2025) 35.79 32.86 34.26 6.32 27.92 10.32 23.41 29.87 26.24 23.08
Table 3: BASIC-L scores across datasets. This metric evaluates low-level structural correspondence between reconstructed and reference images at four granularities: foreground saliency, semantic consistency, instance separation, and part-level delineation.